

********************************************************************
UARS WEEKLY STATUS REPORT - 12 FEB 1999
********************************************************************
***********
SOLSTICE
***********
SOLSTICE has been under daylight only operations via PMON control since
Monday (99039). Full solar scans were completed during this mode but no
stellar calibration experiments were performed. Normal operations resumed
today (Friday 99043).
***********
SUSIM
***********
SUSIM was turned off on Feb. 7 and powered back on on Feb. 12.
On the remaining days, it successfully gathered its normal solar,
offset, and occultation data without apparent problems.
Reprocessing of the SUSIM data to version 20 continues.
***********
PEM
***********
PEM is active PMON control. Energy processor #6 for the HEPS-1 unit
has been found to not activate correctly. Data affected are from the upper
energy EE channel for Head-2 of HEPS-1 (that is the +15 deg electrons above
300 keV). Tests are being conducted on this energy processor to assure
that it is operating correctly. Measurements from this sensor will be
activated as soon as possible. The PEM team and FOT are currently examing
the PMON activation sequence to determine if there is interference during
the commanding of this energy processor.
While under PMON control, the PEM instrument computes energy deposition
from the HEPS unit.
***********
HALOE
***********
HALOE was powered on after the full graze on February 8, 1999. All
engineering parameters are nominal and the science data products are of
excellent quality.
***********
MLS
***********
MLS was reactivated successfully on Jan 26 and operated though Feb 11. The MLS
instrument was operated in a mode emphasizing scientific observations at the
northern latitudes.
***********
HRDI
***********
Nothing received.
***********
WINDII
***********
The testing phase for the WINDII Version 5 processing software
took much longer than expected, but the effort resulted in the
resolution of some long-standing problems with respect to
O2 winds, and Doppler temperatures. This version is now
being promoted to production status.
***********
MPG
***********
Routine processing of instrument Daily Activity Plans continues
with no problems. The MPG continues to assist the Project
Scientist with the scheduling of instrument operations.
***********
CDHF
***********
Building 23 at Goddard, where the CDHF facility resides, has had unexpected
power problems Thursday and Friday. The CDHF is currently up, and was up for
a short time Thursday night (several NRT passes processed). As of this
Friday, we have not been assured that the power problems have been resolved
or that we may still need to power down again.
The CDHF processed current day data through level 3 for all active
instruments. Additionally, CDHF personnel reprocessed: (R)= Requested
(C) = Complete, (F)= Failed (L) = Days remaining to reprocess.
1 day CLA (7.0) ENG (L1-L3) (1R-1C=0L)
0 days HAL (4.80) PROD (L3) (1874R-1721C-1F=152L)
50 days PEM (3.32) PROD (OA-L2) (760R-50C=710L)
197 days SUS (7.00) PROD (L3) (2053R-2042C-11F=0L)
0 days WIN (5.11) ENG (L1-L3) (1R-1F=1L)
2 days SOL (4.62) PROD (TEST) (2R-2C=0L)
CDHF pending software upgrades: DEC AXP Fortran 6.5, DEC AXP OPEN
VMS 6.2, OpenIngres
***********
FOT
***********
REPORTING PERIOD
This report covers 30 January 1999 (Orbit 40362, GMT Day 030) through 05
February 1999 (Orbit 40466, GMT Day 036).
SPACECRAFT OPERATION
The observatory is now in Reverse Flight (Northern Hemisphere viewing) and
is performing nominally under two-battery operations. The instrument
operational changes were:
INSTRUMENT OPERATIONAL CHANGES
Instrument Time Orbit Comment
-No instrument change this report period
The beta angle ranged between the angles of 69.7 degrees to 65.1 degrees
for this report period. The beta angle is now decreasing toward a minimum
angle of -2 degrees for the next yaw on 20 February 1999 (DOY 051).
Spacecraft battery 1, 2, and 3 performance monitoring continues. Battery 1
remains OFF the charge relay. Battery 2 and Battery 3 are currently
maintaining greater than 23.8 V end-of-night (EON) load bus voltage.
Battery temperatures are stable with a temperature delta between Battery 2
and Battery 3 of 0.9 to 1.1 degrees C. Additional battery performance data
for each battery on 29 January 1999 and 05 February 1999 are provided in
Appendix D.
The clock error ranged from 13.1 to -11.7 msec during this report period.
Clock rate adjustments are listed in Appendix C.
All instrument and subsystem engineers were advised on the initial
excursion of ALERTS or Out-Of-Limits conditions detected by the FOT
for this report. A summary is listed below and the details of ALERT and
Out-Of-Limits occurrences are listed in Appendix B.
% ALERTS = 01
% OUT-OF-LIMITS = 05
The following real-time command sheets were executed for the listed
operational element(s) during this report period (see Appendix A).
% FOT 02
% HRDI 02
% PEM 02
UARS Anomaly Reports initiated during this report period are listed in
Appendix E and summarized below.
% FOT 03
One (1) Anomaly Report remained OPEN at the end of this report period (see
Appendix E).
TRANSPONDER FREQUENCIES:
Transponder A: 99/019 - 2287.496951
Transponder B: 99/020 - 2287.499045
GROUND SYSTEM OPERATION
Three (3) operational support problems occurred this period. Anomaly
Reports are included in Appendix E.
The data loss calculations supplied by Data Capture Facility (DCF) from
mission start through UTC Day 99/035 show the observatory data loss to be
21 hours, 50 minutes, 51.540 seconds (an increase of 1.248 seconds since
last report period). This is a 0.0338 percent data loss which equals a
99.9662 percent data capture for the mission.
The increase of 1.248 seconds occurred over two (2) days:
7 0.704 seconds on Day 030 due to insufficient overlap, and
7 0.544 seconds on Day 035 due to data hit.
SOLSTICE operations through the Mission Planning Group (MPG) requested a
set of test 8-week planning aids that start in 1999 and cross into year
2000. The FOT received a partial set of these data this week. These
planning aids will be processed by the FOT next week, after which time, the
MPG will copy the processed files to the UMPGAL system, making them
accessible to the instrument teams.
FOT Operational Database #069 was officially released to the library this
week.
The UARS Test and Training Simulator (UTTS) remains non-operational. The
estimated time to repair the UTTS and return it to operation is unknown.
OTHER SIGNIFICANT EVENTS
UARS SATELLITE OPERATIONS
The FOT continues to support the PEM team with diagnosis of the Energy
Processor 6 status error.
The FOT continues to investigate the feasibility of a new PMON action to
change battery V/T curves based on loading to mitigate battery risks in sun
pointing safehold.
YEAR 2000 OPERABILITY SUPPORT - No change
UARS PROPULSION MODULE B/D THRUSTER LINE FILL
Planning continues for the simulator testing.
UARS RE-ENGINEERING - TELEMETRY PROCESSING - No change
MEETINGS
None.
FUTURE SIGNIFICANT EVENTS
The next UARS Yaw Around Maneuver is scheduled for 20 February 1999.
The next annular solar eclipse will occur on 16 February 1999 over the
Indian Ocean, Australia and the South Pacific Ocean.
PERSONNEL CHANGES
None
Attachments: Appendix A, Real-time Commands Sheets
Appendix B, Out-Of-Limits & ALERT Occurrences
Appendix C, Clock & Solar Array Rate Corrections
Appendix D, UARS Battery Performance Data
Appendix E, UARS Anomaly Reports
APPENDIX A
REAL-TIME COMMAND SHEETS
Week ending 05 February 1999
ORBIT AOS NO. SUBSYS REASON
----- --- --- ------ -------
40368 030/1139 99-046 HRDI Safehold/HI Barrel Temp
40405 032/1956 99-047 PEM Update PDBUs
40406 032/2205 99-048 HRDI Extra daily MicroUs
40416 033/1537 99-049 PEM Update day only sequences
40420 033/2218 99-050 HRDI Load new coefficients
40432 034/1714 99-051 OBC Clk Adj 41.2 to 20.6 msec/day
40455 036/0633 99-052 OBC Clk Adj 20.6 to 41.2 msec/day
40461 036/1539 99-053 HRDI Safehold/HI Barrel Temp
_______________________________________________________________
APPENDIX B
OUT-OF-LIMITS & ALERTS OCCURRENCES
Week ending 05 February 1999
MNEMONIC STATE DESCRIPTION ORBITS
--------- ----- ------------ ------
HRPTSBARTMP Y-LO PTS Barrel Temp 40365-68
HRPTSBARTMP Y-LO PTS Barrel Temp 40397-410
HRPTSBARTMP Y-LO PTS Barrel Temp 40427,28
HRPTSBARTMP Y-LO PTS Barrel Temp 40461
WIQBCURRENT Y-LO WINDII Quiet Bus Current Continuous
The following ALERT occurrences were detected:
MNEMONIC DESCRIPTION ORBITS
--------- ------------ ------
WISRCLSR WINDII Laser Operating (Calibrations) 40428,29
_______________________________________________________________
APPENDIX C
CLOCK & SOLAR ARRAY RATE CORRECTIONS
Week ending 05 February 1999
CLOCK DRIFT RATE ADJUSTMENTS
ORBIT TIME CLOCK RATE CHANGE
----- ---- -----------------
40432 034/17:24 (03 Feb 99) from 41.2 to 20.6 msec/day
40455 036/06:36 (05 Feb 99) from 20.6 to 41.2 msec/day
_______________________________________________________________
APPENDIX D
UARS BATTERY PERFORMANCE DATA
Week ending 05 February 1999
29 Jan 1999 - GMT Day 029 Beta = 69.7 deg, SA Pos = 269 deg
DIFF V (mV) TEMP (C) CURR (amp) EON V AVE C/D AVE
BAT MAX/MIN MAX/MIN MAX/MIN MIN RATIO DOD(%)
--- ------------ ---------- ----------- ------- ------- ------
1 +22.4/ 0.0 -3.68/-4.94 +0.0/ +0.0 16.8 0.000 0.0
2 +22.4/-16.8 +3.87/+2.60 +20.8/-16.0 26.6 1.023 18.1
3 +39.2/-11.2 +2.92/+1.34 +20.0/-15.2 26.6 1.027 17.6
05 Feb 1999 - GMT Day 036 Beta = 65.1 deg, SA Pos = 269 deg
DIFF V (mV) TEMP (C) CURR (amp) EON V AVE C/D AVE
BAT MAX/MIN MAX/MIN MAX/MIN MIN RATIO DOD(%)
--- ------------ ---------- ----------- ------- ------- ------
1 +11.2/-5.6 -3.68/-4.94 +0.0/ +0.0 16.8 0.000 0.0
2 +22.4/-11.2 +3.87/+2.92 +22.8/-16.0 26.4 1.027 18.9
3 +56.0/-11.2 +2.92/+1.97 +22.4/-15.2 26.2 1.030 18.5
V/T LEVEL CHANGES
ORB# TIME LVL FR-TO
----- -------- ----------
no changes V/T 5
APPENDIX E
UARS ANOMALY REPORTS
Week ending 05 February 1999
NEW ANOMALY REPORTS GENERATED
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATALOSS
------ ------ ----- ---- --- ----------------
99-001 FOT 40381 031/0827 14170 Generic Late Acq
99-001 FOT 40385 031/1445 14170 Generic Late Acq
99-006 FOT 40409 033/0427 14170 Generic Late Acq
ANOMALY CLOSURE INFORMATION RECEIVED
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- --- --- -----------------
99-001 FOT see list Jan 98 14170 Generic Late Acq
ANOMALY INVESTIGATION REPORTS GENERATED
AIR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
----- ------ ----- ---- --- -----------------
None this report period
OPEN ANOMALY REPORTS
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
99-006 FOT see list Feb 98 14170 Generic Late Acq
SEVNTFW99/06